# Adapting Differentially Private Synthetic Data to Relational Databases

### v1

@@ 3 Main Results
$\frac{\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{\text{syn}}-\bm{a}\right\|_{2}^{2}}{|\mathcal{W}_{\text{cross},k}|}\leq 8\frac{\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}^{*}-\bm{a}\right\|_{2}^{2}}{|\mathcal{W}_{\text{cross},k}|}+O\left(\frac{d_{\text{max}}}{m\sqrt{\rho_{\text{rel}}}}\right).$
@@
where $\rho_{\text{rel}}=\left(\sqrt{\varepsilon_{\text{rel}}+\operatorname{log}\frac{1}{\delta_{\text{rel}}}}-\sqrt{\operatorname{log}\frac{1}{\delta_{\text{rel}}}}\right)^{2}$ and we ignore logarithmic terms.
@@
The theorem above bounds the per-workload squared error in estimating cross-table queries using synthetic data from Algorithm 1. This upper bound consists of the optimal error without privacy constraints and an additional error term. The optimal error depends on the quality of the individual synthetic tables and the selected privacy mechanism. If the synthetic tables matched the original, the optimal error would be zero. The additional error term is influenced by three factors: $m$ (number of relationships in the original database), $d_{\max}$ (maximum relationships per record), and $\rho_{\text{rel}}$ (privacy parameter). With a fixed $d_{\max}$ and $\rho_{\text{rel}}$, as the number of relationships increases, this error term diminishes. Unlike existing utility theorems that depend on the data domain size $|\mathcal{X}_{i}|$ for categorical features, which can be large, we avoid this reliance in our theorem by using the “marginal trick” [see Appendix D.4 in 27].
4 Conclusion and Limitations
@@
In this paper, we investigated synthetic relational database generation with DP guarantees. We proposed an iterative algorithm that combines with preexisting single-table generation mechanisms to maintain cross-table statistical properties and referential integrity and derived rigorous utility guarantees for it. We hope our efforts inspire new research and push the frontiers of DP synthetic data toward practical scenarios.
@@
There are several intriguing directions worth exploring further. First, we assumed that every record in the relational database must be kept private, but exploring cases where only specific tables contain personal private information would be interesting. For instance, in education data, teacher and student information might require privacy-preserving, while course and department information is often publicly available. Second, our algorithm generates individual synthetic tables and uses an iterative algorithm to establish relationships. An extension would be integrating synthetic table generation into the iterative algorithm, learning both single-table and cross-table queries simultaneously. This may require white-box access to the generative models. Lastly, while we generate synthetic relational databases focused on preserving low-order marginal distributions, other criteria like logical consistency, temporal dynamics, and user-defined constraints are also worthy of exploration.
Appendix A Background on Differential Privacy (DP)
@@
We recall the concept of (zero) concentrated differential privacy [9, 43]. It will be used in the proof of Theorem 1 for establishing DP guarantees for our algorithm.
Definition 3.
@@
$\mathbb{D}_{\alpha}\left(\mathcal{M}(\mathcal{B}),\mathcal{M}\left(\mathcal{B}\right)\right)\leq\rho\alpha$
@@
Lemma 4.
(Relation between zCDP and DP.) If $\mathcal{M}$ satisfies $\rho$-zCDP, then it also satisfies $(\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP for any $\delta>0$.
@@
$\Pr(\mathcal{M}(\mathcal{B})=o)\propto\exp\left(\frac{\epsilon}{2\Delta(\mathsf{score})}\mathsf{score}(\mathcal{B},o)\right),\quad\forall o\in\mathcal{O}.$ (8)
@@
Appendix B More Details about Algorithm 1
We provide more details about our main algorithm (Algorithm 1). Specifically, we include our privacy budget tracking, hyper-parameters, the updating rule for mirror descent, and discretization strategy, in Algorithm 2.
Algorithm 2 Recall Algorithm 1.
Input: private relational database $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$; privacy budget $(\varepsilon_{\text{rel}},\delta_{\text{rel}})$; individual synthetic tables $\mathcal{D}_{1}^{\text{syn}}$, $\mathcal{D}_{2}^{\text{syn}}$; workloads per iteration $K$; number of iterations $T$; hyper-parameter balancing Gaussian and exponential mechanisms $\alpha$; number of relationship in synthetic tables $m^{\text{syn}}$
Convert $(\varepsilon_{\text{rel}},\delta_{\text{rel}})$-DP into $\rho_{\text{rel}}$-zCDP and compute a privacy parameter $\epsilon_{0}=\sqrt{\frac{2\rho_{\text{rel}}}{KT}}$

### v2

@@ 5 Unbiased Sampling
$\left|\frac{1}{m^{\text{syn}}}\bm{q}^{T}\operatorname{vec}(\bm{B}^{\text{syn}})-\frac{1}{m^{\text{syn}}}\bm{q}^{T}\operatorname{vec}(\widetilde{\bm{B}}^{\text{syn}})\right|\leq\frac{1}{m^{\text{syn}}}\cdot\sqrt{\frac{N}{2}\cdot\log\frac{2|\mathcal{Q}|}{\beta}},$
where $N=n^{\text{syn}}_{1}\cdot n^{\text{syn}}_{2}$.
Proof  See Appendix D.6.
The results in the above theorem apply not only to $k$-way marginal queries but also to any statistical queries that can be expressed as a linear function of the bi-adjacency matrix. Additionally, we remark that Srinivasan (2001) also studied the problem of sampling $N$ indicator random variables with a sum constrained to $m^{\text{syn}}$ and proposed an unbiased sampling method. Although their algorithm and ours both have a worst-case time complexity of $O(N)$, empirical comparisons show that our algorithm runs significantly faster when the ratio $m^{\text{syn}}/N$ is small. This improvement arises because, in our Algorithm 3, only the initial merging step requires $O(N)$ operations; thereafter, the probability vector’s dimensionality reduces to $\leq 2m^{\text{syn}}+1$, yielding $O(m^{\text{syn}})$ operations in subsequent steps (see Appendix D.5 for details). In contrast, the runtime of the algorithm in Srinivasan (2001) does not decrease with $m^{\text{syn}}$. In practice, $m^{\text{syn}}/N$ is often quite small (e.g., approximately $4.2\times 10^{-4}$ in the MovieLens dataset), making our algorithm significantly faster than that of Srinivasan (2001).
6 Extension for Improving Scalability
In the previous sections, we presented a projected gradient descent algorithm for learning a relaxed bi-adjacency matrix, along with an unbiased sampling algorithm to convert this matrix back to discrete space. However, the runtime and memory requirements of both algorithms inherently depend on $n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}$, which is the dimension of the bi-adjacency matrix $\bm{B}^{\text{syn}}$ for the synthetic database. In practice, this dimension can be extremely large. For example, in our experiments with the MovieLens dataset (Harper and Konstan, 2015), the user and movie tables contain $6,040$ and $3,900$ records, respectively. To generate a synthetic relational database with the same number of entries, $n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}$ would be nearly $24$ million. This high dimensionality makes it impractical to learn a bi-adjacency matrix that accurately captures the noisy observations derived from real data. To address this challenge, in this section, we introduce a heuristic—random slicing—that can facilitate the synthesis of large-scale relational databases (see Figure 2 for an illustration and Appendix B for its integration into our main algorithm).
The key idea of our heuristic is to explore the sparsity of the bi-adjacency matrix $\bm{B}^{\text{syn}}$. Specifically, we observe that the non-zero entries of $\bm{B}^{\text{syn}}$ typically scale as $O(n_{1}^{\text{syn}}+n_{2}^{\text{syn}})$. For example, in the MovieLens dataset, users tend to rate only a small subset of movies, and in the student-teacher scenario, each student generally enrolls in at most $10$ courses, each taught by a different teacher. In such cases, the number of non-zero entries in $\bm{B}^{\text{syn}}$ (denoted $m^{\text{syn}}$) increases linearly with the size of the user (or student) synthetic table. Based on this, we store $\bm{B}^{\text{syn}}$ as a sparse matrix and update a subset of its elements at each iteration of Algorithm 1.
Figure 2: Illustration of how the random slicing heuristic enhances the scalability of the main algorithm. A full version of the main algorithm, incorporating the random slicing heuristic, is provided in Appendix B.
Our heuristic works as follows. At each iteration of Algorithm 1, we randomly select a fraction $\tau\in(0,1)$ of records from each synthetic table and slice $\bm{B}^{\text{syn}}$ to include only the relationships among the selected records in each table. We denote the sliced bi-adjacency matrix as $\bm{B}^{\text{syn}}_{\text{slice}}$. We then solve a similar relaxed optimization using the projected gradient descent algorithm (Section 4) to optimize the relationships among these selected records, followed by applying the unbiased sampling algorithm (Section 5) to discretize the sliced bi-adjacency matrix. Finally, we re-insert these newly learned relationships into $\bm{B}^{\text{syn}}$. This approach reduces the matrix size for the projected gradient descent and the unbiased sampling algorithms from $n_{1}^{\text{syn}}n_{2}^{\text{syn}}$ to $\tau^{2}n_{1}^{\text{syn}}n_{2}^{\text{syn}}$. Since these algorithms have linear time and space complexity w.r.t. the matrix size, their complexity is reduced to $O(\tau^{2}n_{1}^{\text{syn}}n_{2}^{\text{syn}})$, leading to runtime and memory savings.
7 Utility Guarantees
We present a utility theorem that provides an upper bound on the approximation error of cross-table queries using synthetic relational data generated by Algorithm 1. We begin by defining how to quantify approximation errors for a synthetic relational database $(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{\text{syn}})$. Suppose this synthetic database has $m^{\text{syn}}$ relationships. We stack the query vectors of interest as a matrix $\hat{\bm{Q}}$. Lemma 1 shows that the query answers derived from this synthetic database can be expressed as $\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\operatorname{vec}(\bm{B}^{\text{syn}})$. We now introduce the following definition to measure how accurately $k$-way cross-table queries computed from the synthetic database approximate those obtained from the original data. In this definition, we allow $\bm{B}^{\text{syn}}$ to be a weighted bi-adjacency matrix.
Definition 3
Let $(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{\text{syn}})$ be a synthetic relational database such that $m^{\text{syn}}=\bm{1}^{T}\operatorname{vec}(\bm{B}^{\text{syn}})$. We stack all the $k$-way cross-table query vectors as $\hat{\bm{Q}}$. Denote by $\bm{a}$ the true answers to these queries computed from the original database. We define the (per-workload) mean squared error as
$\mathsf{mse}(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{\text{syn}})\triangleq\frac{1}{|\mathcal{W}_{\text{cross},k}|}\cdot\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\operatorname{vec}(\bm{B}^{\text{syn}})-\bm{a}\right\|_{2}^{2},$ (10)
where $|\mathcal{W}_{\text{cross},k}|$ denotes the total number of $k$-way cross-table marginal workloads.
We now present our utility theorem. It provides an upper bound for the mean squared error in approximating cross-table queries based on outputs from our algorithm. This bound depends on the error associated with the optimal bi-adjacency matrix for the given synthetic tables (i.e., the best achievable error without privacy constraints) and an additional term that approaches zero as the size of the original database increases. The proof techniques are based on classical tools from convex geometry and concentration inequalities.
Theorem 6
Consider running a single iteration of Algorithm 1 (without unbiased sampling) to establish relationships between any individual synthetic tables $\mathcal{D}_{1}^{\text{syn}}$ and $\mathcal{D}_{2}^{\text{syn}}$. The output is a bi-adjacency matrix $\bm{B}^{\text{syn}}$. Let $\bm{B}^{*}$ denote the optimal weighted bi-adjacency matrix in minimizing the mean squared error (Definition 3) when no privacy constraints are enforced. With high probability, we have
$\mathsf{mse}(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{\text{syn}})\leq 8\cdot\mathsf{mse}(\mathcal{D}_{1}^{\text{syn}},\mathcal{D}_{2}^{\text{syn}},\bm{B}^{*})+\widetilde{O}\left(\frac{d_{\text{max}}}{m\sqrt{\rho_{\text{rel}}}}\right),$
@@ 7 Utility Guarantees
where $\rho_{\text{rel}}=\left(\sqrt{\epsilon_{\text{rel}}+\operatorname{log}\frac{1}{\delta_{\text{rel}}}}-\sqrt{\operatorname{log}\frac{1}{\delta_{\text{rel}}}}\right)^{2}$, $m$ is the number of relationships in the original database, and $\widetilde{O}(\cdot)$ ignores logarithmic factors.
Proof  See Appendix D.7.
@@
The theorem above establishes an upper bound on the (per-workload) squared error when estimating cross-table queries using synthetic data generated from Algorithm 1. This upper bound consists of the optimal error without privacy constraints plus an additional error term. The optimal error only hinges on the quality of the individual synthetic tables, determined by the privacy mechanism used for generating these tables. In a hypothetical scenario where the synthetic tables are identical to the original tables, this term would be reduced to zero. On the other hand, the additional error term depends on three factors: $m$ (the number of relationships in the original database), $d_{\max}$ (the maximum number of relationships a record in the original database can have), and $\rho_{\text{rel}}$ (privacy parameter). With fixed $d_{\max}$ and $\rho_{\text{rel}}$, as the number of relationships increases, the additional error term can be reduced to zero. Finally, we note that existing utility theorems (e.g., Liu et al., 2023b; Aydore et al., 2021) often rely on the size of data domain $|\mathcal{X}_{i}|$—that is, the number of distinct values for each categorical feature. This quantity could be large, particularly if data are discretized into numerous bins. We eliminate this dependence in our utility theorem by leveraging the “marginal trick” (see Appendix D.4 in Liu et al., 2021b) in our algorithm.
8 Numerical Experiments
We demonstrate our algorithm through numerical experiments on real-world datasets. The code for reproducing our results is available at https://github.com/azizanlab/dp-relational
Data.
We consider the MovieLens dataset (Harper and Konstan, 2015) and the IPUMS dataset (Ruggles et al., 2021). The MovieLens dataset consists of three tables: users, movies, and ratings. The user table contains 6,040 entries, detailing demographic information like gender, age, and occupation. The movie table contains 3,883 entries, each tagged with associated genres. We convert the genres into 18 binary features—this is not one-hot encoding, as each movie can belong to multiple genres. The ratings table links users and movies in a many-to-many structure: users can rate multiple movies, and each movie can have multiple ratings. We pre-process the rating table to ensure that each movie is rated by no more than 10 users, and each user rates no more than 10 movies (i.e., $d_{\max}=10$). After this pre-processing, the rating table contains 10,075 rows.
The IPUMS dataset contains anonymized U.S. census data. For our experiment, we focus on the user’s table, which provides detailed individual-level data, including demographic, educational, and employment attributes containing 3,373,378 entries. To expedite the training, we only consider the first 5% (168,669 rows) of the records. To establish a many-to-many relationship within this dataset, we use parental links, as the dataset includes the person IDs of each individual’s parents. An individual in the IPUMS dataset may have up to four recorded parents—two mothers and two fathers—leading us to set $d_{\max}=4$. This design creates a self-referential structure within the users table, as the relational connections are entirely within the same dataset table. For the IPUMS dataset, we generate a synthetic users table of 10,000 records.
Setup.
We use state-of-the-art DP mechanisms—AIM (McKenna et al., 2022), MST (McKenna et al., 2021)—to generate individual synthetic tables. Each synthetic table is created with a privacy budget of $\epsilon=1$. We use the OpenDP library (SmartNoise, 2023) to implement these DP mechanisms. To build relationships within the synthetic tables, we apply Algorithm 1 under varying privacy budgets. We evaluate its performance by measuring the average error in $3$-way cross-table marginal queries (Definition 2) between synthetic and real data.
In our ablation study, we vary several parameters. By default, our settings are as follows: the algorithm iterates 15 times, using randomized slicing to create sub-tables of 1,000 synthetic records by 1,000 synthetic records. During each iteration, 3 slices are generated for IPUMS, while 8 are generated in MovieLens, with each slice containing at least 200 rows and 200 columns (20% of the slice) that include relational data to ensure a sufficient number of relationships in each subset. Within each iteration, the exponential mechanism, with a weight of $\alpha=0.2$, selects 3 workloads for evaluation. From the queries selected by the exponential mechanism, we choose 8 of them with the highest error for updating each slice using projected gradient descent.
Results.
We demonstrate the performance of our algorithm across various datasets and DP mechanisms in the figures below. Additionally, we conduct an ablation study to assess the impact of different hyper-parameter selections on our results.
Figure 3: We use AIM and MST to produce individual synthetic tables and then apply our algorithm with privacy budget $\epsilon_{\text{rel}}$ to establish their relationships. We illustrate the impact of varying $\epsilon_{\text{rel}}$ on the average $3$-way marginal query error for the MovieLens dataset (Left) and the IPUMS dataset (Right). As expected, the average error decreases as $\epsilon_{\text{rel}}$ increases, due to reduced noise.
  • Privacy budget. Recall that $\epsilon_{\text{rel}}$ is the privacy budget allocated to Algorithm 1 for establishing relationships among synthetic tables. In Figure 3, we illustrate the effect of varying $\epsilon_{\text{rel}}$ on the average $3$-way marginal query error. As $\epsilon_{\text{rel}}$ increases, the average error decreases due to reduced noise in both the selection (exponential mechanism) and observation (Gaussian mechanism) phases. This trend is clearly observed in results for both AIM and MST. Furthermore, higher $\epsilon_{\text{rel}}$ values lead to a reduction not only in the average error but also in its standard deviation, resulting from the lower noise levels. Finally, with a small privacy budget (e.g., $\epsilon=1$ for generating each synthetic table and $\epsilon_{\text{rel}}=2.0$ for establishing relationships), our algorithm is capable of producing a synthetic relational database with high fidelity.
Figure 4: We analyze the impact of varying the number of iterations $T$ in Algorithm 1 on the quality of synthetic data. The figures show a U-shaped curve between average error and $T$. This trend highlights a trade-off: increasing $T$ allows for more workload queries being selected but their answers, computed from real data, become increasingly noisy.
  • Number of iterations. We analyze how varying the number of iterations $T$ in Algorithm 1 influences synthetic data quality, specifically in terms of the average 3-way marginal query error. Figure 4 shows a U-shaped curve between average error and $T$. This effect arises because the privacy budget $\epsilon_{\text{rel}}$ is evenly distributed across $T$ iterations used by the Gaussian and exponential mechanisms. Hence, increasing $T$ allows more workloads to be selected and optimized for synthetic data with the cost of adding higher noise to the workload query answers computed from the real data. Conversely, with very few iterations, the algorithm has limited chances to explore sufficient workload queries, leading to higher average error. This trade-off underpins the U-shaped curve, identifying an optimal range between $T=10$ and $T=15$ where the error is minimized.
Figure 5: We show the effect of the hyperparameter $\alpha\in[0,1]$ on the quality of synthetic data produced by our algorithm. This parameter allocates the privacy budget between the Gaussian and exponential mechanisms.
  • Privacy budget allocation. In Algorithm 4, a detailed version of our algorithm, we introduce a hyperparameter, $\alpha\in[0,1]$, to divide the privacy budget between the Gaussian and exponential mechanisms. When $\alpha=1.0$, the entire privacy budget is allocated to the exponential mechanism, enabling it to select the most relevant workloads. However, with no privacy budget assigned to the Gaussian mechanism, it outputs random values for the selected workload queries. Conversely, setting $\alpha=0.0$ dedicates the entire privacy budget to the Gaussian mechanism, producing low-noise query values derived from real data for randomly selected workloads. Figure 5 illustrates the impact of varying $\alpha$ on the quality of the synthetic data. Based on this result, we select $\alpha=0.2$ for the rest of our ablation study.
Figure 6: We demonstrate the effect of the number of workloads $K$ selected by the exponential mechanism on the quality of synthetic data produced by our algorithm.
  • Number of workloads selected by the exponential mechanism. The exponential mechanism in Algorithm 1 selects $K$ workloads per iteration. With the privacy budget distributed evenly across these selected workloads, a larger $K$ increases the noise added to their query answers computed from real data, while a smaller $K$ may yield an insufficiently diverse set of observed workloads. Figure 6 illustrates the impact of different values of $K$ on the average $3$-way marginal error.
Figure 7: We demonstrate the effect of the number of workloads optimized by the relaxed optimization. As more workloads are included, the average $3$-way marginal query error reduces.
  • Number of workloads optimized by the relaxed optimization. To enhance the speed of our algorithm, we select a fixed number of workloads with the highest error each time we solve the relaxed optimization, choosing from all the workloads identified by the exponential mechanism so far. This selection strategy prioritizes computational efficiency, so we anticipate that including more workloads per step should reduce the average $3$-way marginal query error. Figure 7 validates this hypothesis, showing a decline in error as the number of workloads grows.
9 Conclusion, Future Work, and Impact Statement
@@
In this paper, we investigated synthetic relational database generation with DP guarantees. We proposed an algorithm that can be combined with any preexisting single-table generation mechanisms to maintain cross-table statistical properties and referential integrity. This algorithm creates relationships between individual synthetic tables by learning a bi-adjacency matrix, which is iteratively updated through solving a quadratic program and sampling an unbiased instance to minimize approximation errors of $k$-way marginals. We derived rigorous utility guarantees for our algorithm and provided a PyTorch implementation suited for high-dimensional and high-volume data. Furthermore, we conducted comprehensive experiments and introduced new benchmark datasets and evaluation metrics to assess the performance and scalability of our algorithm. We hope our effort can inspire new research and push the frontiers of DP synthetic data towards more practical scenarios.
@@
There are several intriguing directions that deserve further exploration. First, we assumed any record in the relational database must be kept private. It would be interesting to explore scenarios in which only specific tables contain personal private information. For example, in the context of education data, it may be pertinent to examine situations where only teacher and student information requires privacy-preserving, while course and department information, being often publicly available, does not. Second, our algorithm generates individual synthetic tables and uses an iterative algorithm to establish their relationship; one extension would be integrating synthetic table generation into the iterative algorithm, learning both single-table and cross-table queries simultaneously. However, achieving this may require white-box access to the generative models. Finally, while we generate synthetic relational databases with the goal of preserving low-order marginal distributions, there are other criteria worthy of exploration, such as logical consistency, temporal dynamics, and user-defined constraints. It would be valuable to understand whether these criteria can be incorporated into synthetic relational data generation while maintaining DP guarantees.
This paper focuses on privacy-preserving synthetic data generation, building upon prior efforts in DP synthetic data. We extend their application to a more practical scenario where data is stored in a relational database. This line of research may significantly impact several critical domains, such as finance, healthcare, and government, where safeguarding data privacy is paramount. The deployment of DP synthetic data can lead to more inclusive research practices, allowing organizations to share data without compromising the privacy of individuals’ personal information. This, in turn, facilitates collaboration and propels advancements in research and business applications reliant on data-driven insights.
It is crucial to acknowledge potential challenges and ethical considerations associated with the deployment of synthetic data. Synthetic data proves to be a suitable substitute for tasks that do not necessitate absolute precision, such as data visualization, software testing, and initial model development. Nevertheless, there is often a trade-off between preserving the essential statistical properties of the original data and meeting privacy requirements. For applications with individual-level consequences, any methods or insights derived from synthetic data need to be carefully evaluated to avoid unintended consequences and biases in downstream applications. In summary, while our research advances the field of DP synthetic data generation, we recognize the importance of ethical considerations and the potential societal impact. By mitigating privacy concerns in data sharing, this work aspires to play a pivotal role in fostering the responsible and reliable development of advanced machine learning technologies.
A Background on Differential Privacy (DP)
@@
We recall the concept of (zero) concentrated differential privacy (Bun and Steinke, 2016; Steinke, 2022). It will be used in the proof of Theorem 1 for establishing DP guarantees for our algorithm.
Definition 4
@@
$\mathbb{D}_{\alpha}\left(\mathcal{M}(\mathcal{B}),\mathcal{M}\left(\mathcal{B}\right)\right)\leq\rho\alpha,$
@@
Lemma 4
(Relation between zCDP and DP). If $\mathcal{M}$ satisfies $\rho$-zCDP, then it also satisfies $(\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP for any $\delta>0$.
@@
$\mathbb{P}(\mathcal{M}(\mathcal{B})=o)\propto\exp\left(\frac{\epsilon}{2\Delta(\mathsf{score})}\mathsf{score}(\mathcal{B},o)\right),\quad\forall o\in\mathcal{O}.$ (13)
@@
B More Details about Algorithm 1
In this section, we outline the implementation details of our main algorithm (Algorithm 1). The full version is presented in Algorithm 4 and illustrated in Figure 2. Recall that our algorithm selects worst-case workloads using the exponential mechanism and computes noisy observations of these workloads from real data using the Gaussian mechanism. Details on these mechanisms can be found in Algorithms 5 and 6, along with their respective privacy budget allocations. Additionally, we incorporate the random slicing mechanism (Section 6) to enhance scalability for generating large-scale synthetic relational databases.
To implement the random slicing mechanism, we begin by storing the bi-adjacency matrix as a sparse matrix $\bm{B}^{\text{syn}}\in\{0,1\}^{n_{1}^{\text{syn}}}\times\{0,1\}^{n_{2}^{\text{syn}}}$. We introduce a parameter $\tau<1$, which specifies the fraction of relations to be learned in each iteration. Next, we define the slicing operation, which will be used to extract elements from $\bm{B}^{\text{syn}}$ before running the projected gradient descent and unbiased sampling algorithms.
Definition 5
Given a matrix $\bm{M}\in\mathbb{R}^{n_{1}\times n_{2}}$ and vectors $\bm{v}_{1}\in\{0,1\}^{n_{1}}$, $\bm{v}_{2}\in\{0,1\}^{n_{2}}$, we define the slicing operation as $\left[\bm{v}^{1}\right]\bm{M}\left[\bm{v}^{2}\right]$. This operation yields a matrix with $\bm{1}^{T}{\bm{v}^{\text{slice}}_{1}}$ rows and $\bm{1}^{T}{\bm{v}^{\text{slice}}_{2}}$ columns. The $(i,j)$ entry of this matrix corresponds to the $(v^{1}_{i},v^{2}_{j})$ entry of $\bm{M}$, where $v^{1}_{i}$ is defined as the $i$-th nonzero entry of $\bm{v}^{1}$ (similarly for $v^{2}_{j}$). If only the columns are sliced, we write $\bm{M}\left[\bm{v}^{2}\right]$.
Algorithm 4 Full Algorithm 1 (detailed version with scalability enhancements).
Input: private relational database $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$; privacy budget $(\epsilon_{\text{rel}},\delta_{\text{rel}})$; individual synthetic tables $\mathcal{D}_{1}^{\text{syn}}$, $\mathcal{D}_{2}^{\text{syn}}$; workloads per iteration $K$; number of iterations $T$; number of relationship in synthetic tables $m^{\text{syn}}$; slicing fraction $\tau$; hyper-parameter balancing Gaussian and exponential mechanisms $\alpha$
Initialize: $\hat{\bm{Q}}=\emptyset,\hat{\bm{a}}=\emptyset$, $n_{1}^{\text{syn}}$ by $n_{2}^{\text{syn}}$ sparse matrix $\bm{B}^{\text{syn}}$ with entries $\in\{0,1\}$
Convert $(\epsilon_{\text{rel}},\delta_{\text{rel}})$-DP into $\rho_{\text{rel}}$-zCDP and compute a privacy parameter $\epsilon_{0}=\sqrt{\frac{2\rho_{\text{rel}}}{KT}}$